OpenStudy (anonymous):

Which is the set of roots of the equation? 0=x^2+10x-24 {12, 2} {-12, -2} {-12, 2} {12, -2}

OpenStudy (hoblos):

x^2+10x-24= (x+12)(x-2) => solutions are {-12, 2}

OpenStudy (anonymous):

Factor and solve ...

OpenStudy (anonymous):

Hi!

OpenStudy (anonymous):

Do you know how to factor?

hero (hero):

x^2 + 10x - 24 = 0 x^2 + 12x - 2x - 24 = 0 x(x+12)-2(x+12) =0 (x+12)(x-2) = 0 x = -12, 2